First you are given a table with 2 people on each side and one on each end.
Then you are asked to use y=mx+b to create a formula telling how many people could be seated at 4 tables. Keep in mind each time you add a table you lose one of your end seats because the tables are put end to end.

Answer by richwmiller(17219)

y=6 x=1
y=11 x=2
y=16 x=3
11-6 =m(2-1)
5=m
y=mx+b
6=5*1+b
b=1
y=5x+1